OpenStudy (damonte212):

What is the range of the function f(x) = 12 − 3x for the domain {-4, -2, 0, 2, 4}? Medal/Fan A.{24, 18, 12, 6, 20} B.{6, 12, 18, 24, 30} C.{-12, -6, 0, 6, 12} D.{0, 6, 12, 18, 24} @M3GA-BIT3 @jaceykilla @Og_Lover_smokin_

OpenStudy (anonymous):

Tbh Ii Don Kno

OpenStudy (damonte212):

right this shhh hard asf

OpenStudy (anonymous):

give Me My Medal Lol

OpenStudy (damonte212):

Yeet!!!

OpenStudy (kamibug):

Range is the y values and domain is the x values. Remember "f(x)" is the same thing as "y." So plug in each given domain value and solve for y. For example, for the first domain value, -4, plug it in and solve... f(x) = 12 − 3(-4) f(x) = 12 + 12 f(x) = 0 y = 0 So you know the first y value of your answers has to be 0. So I basically gave you the answer, Lol. Anyways, you'd do that with each x value to find all the other range values. :)

OpenStudy (damonte212):

???

OpenStudy (m3ga-bit3):

lol ya what she said:) imma give u a medal:)

OpenStudy (m3ga-bit3):

@damonte212 the only answer with a 0 in the beginning of it...:XD

OpenStudy (damonte212):

ight

OpenStudy (anonymous):

Ima Weak Asf

OpenStudy (m3ga-bit3):

hey guys fan me plz and if u need help with anything besides math id live to help:) oh and add me on face book Boeck Megan:p

OpenStudy (m3ga-bit3):

y u weak @Og_Lover_smokin_ nvm i dont wanna kno

OpenStudy (damonte212):

add me @M3GA-BIT3 Damonte To Real Moore

OpenStudy (m3ga-bit3):

ur added:)

OpenStudy (anonymous):

@ cazuse Me damonte in the same class and he askin me question and don put my name inna comment df

OpenStudy (m3ga-bit3):

XD

OpenStudy (m3ga-bit3):

my bad

OpenStudy (m3ga-bit3):

some one woke up on the wrong side of the bed